Grade 12th passGeneral Physics

A body radiates 5 W energy at atemperature of 400 K . If the temperature is increased to 1200 K . then it radiates energy at the rate of
1. 405w
2.
3.
4. 45w

G Sir Askiitians faculty
8 Years agoGrade 12th pass

1 Answer

Eshan
8 Years ago
The rate of energy radiation from a body at tempertare T,

P=\sigma A T^4

Hence\dfrac{P_2}{P_1}=\dfrac{T_2^4}{T_1^4}
\implies P_2=3^4\times 5W=405W